the sum of ages of the son and his father is 69. if the father 's age is twice as his son's age increased by 9, find their present ages.
Answer by Maths68(1474)

Let
f= father’s current age
s=son’s current age
the sum of ages of the son and his father is 69.
f+s=69
s=69-f………….(1)
son's age increased by 9 = s+9
Father 's age is twice as his son's age increased by 9 means
f=2(s+9)
f=2s+18…………(2)
Put the value of s from (1) to (2)
f=2(69-f)+18
f=138-2f+18
f=-2f+156
f+2f=156
3f=156
f=156/3
f=52
Put the value of f in (1)
s=69-f
s=69-52
s=17
Father is 52 and Son is 17 years old at present.